OpenStudy (anonymous):

integrate 1/(4+x^2) x=0 , x=(2/squareroot3)

OpenStudy (anonymous):

\[\int\limits_{0}^{2/\sqrt{3}}1/(4+x ^{2})\] =(1/4)*1/(1+(x/2)^2) \[=1/4 *\tan^{-1} x/2\](x=0 to 2/root3) so by putting the value of x u will get the ans as \[\pi/6\]